• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Solvignon

    Log entry time 16:00:02 on March16,2011

    Entry number 352153

    keyword=Day shift summary

    8:00 starting the day shift in production mode on 4He.

    9:19 Hall C is started the cooldown of their target. We moved our target to slanted C.

    9:30 Beam study starts. Ask to have Hall A in controlled access.

    13:00 Beam study done. Waiting for beam.

    13:20 Tuned beam.

    13:28 Done with tuned beam. Moved target from empty to BeO.

    13:35 Asked for a Harp scan. Beam size good in 4A: x=124, y=155. 4B scan cannot be done due to hardware problem.

    13:49 Start pointing and optics study.

    14:25 Moved target from BeO to optics.

    14:28 Ready for beam.

    15:19 Start moving the Left sieve slit in.

    15:25 Left sieve slit is in.


    ----------------------------------------------------------------------
    Run summary:
    ------------
    * kinematics: Left Pe=3.6014 GeV/c, thetae=20.3deg
    Right Pp=1.3825 GeV/c, thetap=33.5deg
    I_BB = 518A



    07:59-08:25 run# 2865 He4 275K events problem with HAND ND N6 left ADC
    08:28-09:20 run# 2866 He4 557K events good
    13:52-14:25 run# 2868 Be0 202K events raster off
    14:31-15:03 run# 2869 optics 181K events 10uA
    15:07-15:16 run# 2870 optics 29K events 10uA (test F1)
    15:27-??-?? run# 2871 optics ongoing 25uA, left sieve in